# A Practical Approach for Exploring Granger Connectivity in High-Dimensional Networks of Time Series

### v1

@@ Abstract
This paper presents a novel method for discovering effective connectivity between specified pairs of nodes in a high-dimensional network of time series. To accurately perform Granger causality analysis from the first node to the second node, it is essential to eliminate the influence of all other nodes within the network. The approach proposed is to create a low-dimensional representation of all other nodes in the network using frequency-domain-based dynamic principal component analysis (spectral DPCA). The resulting scores are subsequently removed from the first and second nodes of interest, thus eliminating the confounding effect of other nodes within the high-dimensional network. To conduct hypothesis testing on Granger causality, we propose a permutation-based causality test. This test enhances the accuracy of our findings when the error structures are non-Gaussian. The approach has been validated in extensive simulation studies, which demonstrate the efficacy of the methodology as a tool for causality analysis in complex time series networks. The proposed methodology has also been demonstrated to be both expedient and viable on real datasets, with particular success observed on multichannel EEG networks.
1 Introduction
Identifying Granger-causal connections between nodes in high-dimensional networks in neuroscience and network science has significant outcomes. This is particularly relevant when analyzing EEG/fMRI data, where the attention is on a few channels/regions of interest within an extensive network of interacting nodes (Friston, 2011; Friston et al., 2013; Yang et al., 2016; Tu et al., 2019). Conventional Granger causality (GC) is the standard approach to infer causal relationships in high-dimensional EEG networks. However, they are subject to substantial limitations due to the impact of extraneous channels on the Channels of Interest (COI). Accordingly, there is a need for an effectual and parsimonious method to isolate and accurately evaluate the COI’s causal interactions, especially in the presence of multiple channels. The prevailing approaches to causality analysis in high-dimensional networks typically depend on F-tests for Granger causality, which are predicated on the assumption of Gaussianity. However, real-world datasets frequently violate this assumption, leading to biased and unreliable results. Moreover, these methodologies need improvement in effectively removing the influence of non-involved nodes on the COI, which would otherwise lead to misrepresenting causal relationships. On the other hand, conventional dimension reduction techniques in the time domain, such as principal component analysis (PCA), cannot fully capture the data’s frequency-specific characteristics, which are crucial for understanding neural dynamics.
@@ 1 Introduction
We propose a practical and easily accessible approach based on spectral dynamic principal component analysis (sDPCA) in the frequency domain, combined with permutation testing for Granger causality. The method sDPCA, initially proposed by Brillinger (1969, 1981), extends static PCA by incorporating temporal dependency within the data, thereby reducing the dimensionality of time series in the frequency domain while retaining maximal information. Various methodologies have evolved from sDPCA. For instance, Stoffer (1999) developed a spectral envelope using sDPCA to identify common signals across multiple time series, while Hörmann et al. (2015) proposed a functional variant of sDPCA for analyzing functional time-series data. Ombao and Ho. (2006), Yuxiao Wang et al. (2019), and Ombao and Pinto (2022) indicated that spectral principal component analysis (sPCA) in the frequency domain offers practical benefits and is highly effective in low-dimensional representations of high-dimensional time series. The main objective of this study is to illustrate how the proposed approach enables the practical and sufficient isolation of COI by partialling out the influence of disinterested nodes, which are derived as a dynamic summary in a low-dimensional representation. The frequency-domain dynamic PCA method maintains the spectral characteristics of the data, which are crucial for accurately determining causal relationships in time series networks. By employing a permutation-based GC test in lieu of the conventional F-test, we can relax with the assumption of normality in the error terms, thereby assuring robustness to deviations from normality and rendering the proposed approach more suitable for real-world data.
@@
The proposed approach offers two main advantages. Firstly, DPCA in the frequency domain offers an accurate and practical dimension reduction while preserving the spectral information. This is also essential because EEGs are correlated across time, and hence, the spectrum captures cross-covariances at all lags. This leads to a more precise isolation of COI, reducing the likelihood of spurious causal inferences. Secondly, the permutation-based GC test improves the reliability of causality detection by not relying on the assumption of Gaussianity. This allows the test to accommodate distributions that deviate significantly from normality, such as the fat-tailed distributions observed often in EEG data. The procedure is straightforward and practical to apply. The subsequent sections will elaborate on the proposed approach and its components. Section 2 introduces the Spectral DPCA method and the permutation GC test. Section 3 outlines the scope of the simulation study and presents related results. Section 4 applies our approach to real EEG data and discusses the findings.
2 Methodology
2.1 Inference in high-dimensional networks of time series
@@
Define $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ to be a high dimensional network of time series with possibly $n\geq 256$ nodes. As we consider in this paper, this network may represent the activity of a brain through EEG recordings in which each component (i.e., node or channel) is actually a time series. In this notation of network $N$, $Y^{*}$ and $X^{*}$ denote any two nodes of interest. For problems whose solutions are sought in component-wise, we set our goal as exploring the interaction or efficient connectivity between any pair of nodes in a network rather than the relationality of the whole network. These two nodes are subject to the influence of other nodes due to the network’s relational structure. This means that to dissect the respective efficient connection between any two nodes of interest, they need to be freed from the influences of other nodes. Thereby, the interference-free counterparts of the nodes of interest can be obtained in the following manner:
${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}Y_{t}=Y^{*}_{t}-E\big{(}Y^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}X_{t}=X^{*}_{t}-E\big{(}X^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
@@
where the $f_{z}(.)$ is a function that ensures the dynamic variability of excluded nodes. Practically, the function $f_{z}(.)$ acts as a filter and can take any form as long as it provides the net summary discharge of the nodes excluded. Within the scope of this study, we utilize the spectral domain dynamic PCA implementation made available by Hörmann et al. (2015) via R package called freqdom. This method, originally proposed by Brillinger (1981), enables us to obtain principal component scores that concisely encapsulate variability in multi-dimensional time series, specifically with regard to their serial dependence (see, e.g., Shumway and Stoffer (2017), Ombao and Pinto (2022)). Hence, we propose the filter $f_{z}(.)$ over excluded nodes to provide the following summary indicators,
@@
where $pc1,pc2,\ldots,pc_{k}$ are the scores of the principal components, and interactions represent any interaction terms between these principal components.
@@
$E\big{(}Y^{*}_{t}\mid f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)}=a_{1}\cdot pc1+a_{2}\cdot pc2+\cdots+a_{k}\cdot pc_{k}+\sum a_{ij}\cdot(pc_{i}\times pc_{j}),$
$E\big{(}X^{*}_{t}\mid f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)}=b_{1}\cdot pc1+b_{2}\cdot pc2+\cdots+b_{k}\cdot pc_{k}+\sum b_{ij}\cdot(pc_{i}\times pc_{j}),$
@@
$\hat{C}_{\mathbf{Z}}(h)=\frac{1}{T}\sum_{t=|h|+1}^{T}(\mathbf{Z}_{t+|h|}-\bar{\mathbf{Z}})(\mathbf{Z}_{t}-\bar{\mathbf{Z}})^{\prime},$
@@
2.2 Permutation-based Granger Causality test
Granger Causality (GC) is a statistical hypothesis test used to determine whether one time series can predict another. The concept, introduced by Granger (1969), is based on the principle that if the prediction of a time series $Y_{t}$ can be improved by incorporating past values of another time series $X_{t}$, then $X_{t}$ is said to "Granger-cause" $Y_{t}$.
@@ 2.2 Permutation-based Granger Causality test
Permutation-based Granger causality (GC) testing offers a robust alternative to conventional GC tests by relaxing the Gaussianity assumption. The permutation-based GC test evaluates the significance of causality by comparing the observed test statistic against a distribution of test statistics generated through random permutations of the data. This approach is more robust to deviations from normality and other potential violations of the assumptions underlying conventional GC.
To perform the permutation test, we first calculate the test statistic, $F$, for the observed data. We then generate a distribution of the test statistic under the null hypothesis by randomly permuting the time series $X_{t}$ and recomputing the F-statistic for each permutation. Specifically, we shuffle $X_{t}$ to break any potential temporal structure and repeat the GC test for each permuted dataset. This creates a distribution of F-statistics that represent the null distribution.
The permutation p-value is then computed as the proportion of permuted test statistics that are greater than or equal to the observed test statistic:
$p_{\text{perm}}=\frac{\sum_{k=1}^{N}\mathbf{1}(F_{k}\geq F_{\text{obs}})}{N}$
where $N$ is the number of permutations, $F_{k}$ is the test statistic for the $k$-th permutation, $F_{\text{obs}}$ is the observed test statistic, and $\mathbf{1}(\cdot)$ is the indicator function. The indicator function $\mathbf{1}(F_{k}\geq F_{\text{obs}})$ takes the value 1 if $F_{k}$ is greater than or equal to $F_{\text{obs}}$ and 0 otherwise. This function effectively counts the number of times the permuted test statistic exceeds the observed test statistic, providing a measure of how extreme the observed value is within the context of the null distribution.
This permutation approach provides a non-parametric method to assess the significance of the causality, making it more flexible and robust compared to traditional methods. By comparing the observed F-statistic to the distribution generated from permuted data, we account for the actual data distribution and potential dependencies, leading to more reliable inference.
3 Simulation study
@@
The aim of the simulation study is to evaluate the effectiveness of spectral dynamic principal component analysis (sDPCA) combined with Granger causality (GC) testing in identifying true causal relationships within a high-dimensional network of time series data, specifically focusing on EEG data. The study investigates how well the proposed method isolates the channels of interest (COI) from the influences of other channels and accurately assesses their causal interactions.
The study begins by generating a set of 10 channels as mixtures of latent processes proposed by Ombao and Pinto (2022). Each latent process is modeled as an autoregressive process of order 2 (AR(2)). Mathematically, the channels and corresponding indexes are represented as:
$Z_{i}(t)=a_{i}\cdot\delta(t)+b_{i}\cdot\theta(t)+c_{i}\cdot\alpha(t)+d_{i}\cdot\beta(t)+e_{i}\cdot\gamma(t)+\epsilon_{i}(t)$
@@
for $i=1,\ldots,5,11,\ldots,15$, where $\delta,\theta,\alpha,\beta,\gamma$ are latent AR(2) processes, and $\epsilon_{i}(t)$ are error terms.
Next, we generate another set of 10 channels, where each channel is caused by a corresponding channel from the first set with added noise. This can be represented as:
$Z_{i+5}(t)=\phi_{1}Z_{i}(t-1)+\phi_{2}Z_{i}(t-2)+\phi_{3}Z_{i}(t-3)+\phi_{4}Z_{i}(t-4)+\epsilon_{i+5}(t)$
for $i=1,\ldots,5,11,\ldots,15$. The error terms $\epsilon_{i+5}(t)$ follow a specified distribution (e.g., t-distribution with degrees of freedom 15, normal distribution, or non-normal distribution with specified skewness and kurtosis).
The ground truth for Granger causality in this setup is defined by the following relationships:
$Z_{1}\to Z_{6},\quad Z_{2}\to Z_{7},\quad Z_{3}\to Z_{8},\quad Z_{4}\to Z_{9},\quad Z_{5}\to Z_{10},\quad Z_{11}\to Z_{16},\quad Z_{12}\to Z_{17},\quad Z_{13}\to Z_{18},\quad Z_{14}\to Z_{19},\quad Z_{15}\to Z_{20}$
The network is then extended by generating an additional 236 channels ($Z_{21}$ to $Z_{256}$) that share common latent processes with the first 20 channels. This ensures connectivity within the network. Some of these new channels are allowed to have linear influences on the first 20 channels and can be defined as:
$Z_{i}(t)=Z_{i}(t)+\sum_{j\in\text{influencing set}}\text{weight}\cdot Z_{j}(t)$
The number of influencing channels and the weight of their influence are varied in different scenarios to study their effect on the performance of the GC tests.
The study then applies sDPCA to the extended network, focusing on the 236 additional channels to isolate the COI (the first 20 channels). The principal component scores are used to remove the effects of the other channels on the COI. The adjusted time series are obtained by regressing the COI on the principal component scores and retaining the residuals. Bivariate component-wise VAR(1) models are fitted to the residuals to compute the F-statistics for the GC tests. These F-statistics are then compared to critical values from the F-distribution to decide on the presence of GC.
Finally, permutation-based F-statistics are computed by repeatedly permuting the predictor time series and recalculating the F-statistics. The proportion of permuted F-statistics that exceed the observed F-statistic provides the permutation p-value, offering a robust measure of causality that accounts for the actual data distribution.

### v2

# Granger Causality in High-Dimensional Networks of Time Series

@@ Abstract
A novel approach is developed for discovering directed connectivity between specified pairs of nodes in a high-dimensional network (HDN) of brain signals. To accurately identify causal connectivity for such specified objectives, it is necessary to properly address the influence of all other nodes within the network. The proposed procedure herein starts with the estimation of a low-dimensional representation of the other nodes in the network utilizing (frequency-domain-based) spectral dynamic principal component analysis (sDPCA). The resulting scores can then be removed from the nodes of interest, thus eliminating the confounding effect of other nodes within the HDN. Accordingly, causal interactions can be dissected between nodes that are isolated from the effects of the network. Extensive simulations have demonstrated the effectiveness of this approach as a tool for causality analysis in complex time series networks. The proposed methodology has also been shown to be applicable to multichannel EEG networks.
Index Terms:
High-Dimensional Networks of Signals, Connectivity, Causality, Spectral Dynamic PCA, EEG signals
I Introduction
Identifying effective connectivity in high-dimensional networks (HDN) is a challenging problem that remains an active area of research and has been previously addressed by several studies [see, e.g., 1, 2, 3, 4, 5, 6, 7, 8]. Moreover, it is difficult to infer or interpret interactions between a pair of nodes while taking into account potential interactions with all other nodes in the background [9, 10]. This is a particularly relevant issue (e.g., analysis of EEG-like signals) when the focus might stand on a few channels/nodes of interest (COI/NOI) within an extensive network of interacting nodes. In fact, utilizing conventional vector autoregressive models (VAR) can be a practical way to identify pairwise causal interactions in high-dimensional networks as long as the potential confounding effect due to the high-dimensional background nodes is addressed. In this paper, we propose a framework that provides a low-dimensional summary of the background activity in the high-dimensional network (HDN).
To discern effective connectivity, particularly between pairs of nodes in an HDN of brain signals, we propose a practical and easily accessible approach based on spectral dynamic principal component analysis (PCA) ([11, 12]) combined with testing for GC in lieu of the conventional F-test.
The sDPCA method enhances the classical PCA by incorporating temporal dependency within the data, thereby reducing the dimensionality of time series in the frequency domain while retaining maximal temporal information. In other words, while the classical PCA is based on the eigendecomposition of the zero-lag (or contemporaneous) covariance matrix, sDPCA is based on the eigendecomposition of the spectral matrix at all frequencies. Since there is a 1-1 relationship between the spectral matrix (at all frequencies) and the covariance matrix (at all lags) then this approach captures lead-lag relationships amongst signals.
@@ I Introduction
Various methodologies have evolved from sDPCA. For instance, [13] developed a spectral envelope using sDPCA to identify common signals across multiple time series, while [14] proposed a functional variant of sDPCA for analyzing functional time-series data. As indicated by [15, 5, 16], spectral principal component analysis (sPCA) in the frequency domain offers practical benefits and is highly effective in low-dimensional representations of high-dimensional time series. The main objective of this study is to illustrate how the proposed approach enables the practical and sufficient isolation of COI by partialling out the influence of disinterested nodes, which are derived as a dynamic summary in a low-dimensional representation. The frequency-domain dynamic PCA method maintains the spectral characteristics of the data, which are crucial for accurately determining causal relationships in time series networks.
@@
The proposed approach offers two main advantages. Firstly, sDPCA in the frequency domain offers an accurate and practical dimension reduction while preserving the spectral information. This is essential because the spectrum of a signal captures cross-covariances at all lags. This leads to a more precise isolation of COI, reducing the likelihood of spurious causal inferences. Secondly, the proposed approach enables the formation of a comprehensive yet interpretable representation of the interactions between specified nodes at a defined level of resolution rather than being unable to infer by the expansive number of possible interactions between nodes. This facilitates the study of considerable numbers of interactions, ranging from pairs of nodes to interactions between a tractable number of clusters of nodes. The subsequent sections will elaborate on the proposed approach and its components. Section 2 presents the proposed approach. Section 3 outlines the scope of the simulation study and presents related results. Section 4 applies the proposed approach to EEG data and discusses the findings.
II Methodology
II-A Inference in high-dimensional networks of time series
@@
Define $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ to be a high dimensional network of signals with possibly $n\geq 256$ nodes. As we consider in this paper, this network may represent the activity of a brain through EEG recordings in which each component (i.e., node or channel) is actually a time series. In this notation of network $N$, $Y^{*}$ and $X^{*}$ denote any two nodes of interest. For problems whose solutions are sought component-wise, we set our goal as exploring the interaction or efficient connectivity between any pair of nodes in a network rather than the relationality of the whole network. These two nodes are subject to the influence of other nodes due to the network’s relational structure. This means that to dissect the respective efficient connection between any two nodes of interest; they need to be freed from the influences of other nodes. Thereby, the interference-free counterparts of the nodes of interest can be obtained in the following manner:
$Y_{t}=Y^{*}_{t}-E\big{(}Y^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
$X_{t}=X^{*}_{t}-E\big{(}X^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
@@
where the $f_{z}(.)$ is a function that encompasses the dynamic variability of excluded nodes. Practically, the function $f_{z}(.)$ acts as a filter and can take any form as long as it provides the net summary discharge of the nodes excluded. Within the scope of this study, we utilize the spectral domain dynamic PCA implementation made available by Hörmann et al. (2015) via R package called freqdom. This method, originally proposed by Brillinger (1981), enables us to obtain principal component scores that concisely encapsulate variability in multi-dimensional time series, specifically with regard to their serial dependence [17, 16, see, e.g., ]. Hence, we propose the filter $f_{z}(.)$ over excluded nodes to provide the following summary indicators,
@@
where $pc1,pc2,\ldots,pc_{k}$ are the scores of the principal components obtained through sDPCA, and interactions represent any interaction terms between these principal components.
@@
$E\big{(}Y^{*}_{t}\mid f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)}=a_{1}\cdot pc1+a_{2}\cdot pc+$
$\cdots+a_{k}\cdot pc_{k}+\sum a_{ij}\cdot(pc_{i}\times pc_{j}),$
$E\big{(}X^{*}_{t}\mid f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)}=b_{1}\cdot pc1+b_{2}\cdot pc2+$
$\cdots+b_{k}\cdot pc_{k}+\sum b_{ij}\cdot(pc_{i}\times pc_{j}),$
@@
$\hat{C}_{\mathbf{Z}}(h)=\hat{C}_{\mathbf{Z}}(-h)^{\prime}=\frac{1}{T}\sum_{t=1}^{T+h}(\mathbf{Z}_{t}-\bar{\mathbf{Z}})(\mathbf{Z}_{t-h}-\bar{\mathbf{Z}})^{\prime}.$
@@
Terms and Notation
  • $\mathbf{Z}$: Original multivariate time series matrix of dimensions $T\times n$, where $T$ is the number of time points and $n$ is the number of variables.
  • $F_{\mathbf{Z}}(\omega)$: Cross-spectral density matrix at frequency $\omega$.
  • $\text{Cov}(\mathbf{Z}_{h},\mathbf{Z}_{0})$: Covariance matrix at lag $h$.
  • $\hat{C}_{\mathbf{Z}}(h)$: Empirical lagged auto-covariance matrix at lag $h$.
  • $w(\cdot)$: Window function used to smooth the estimates.
  • $q$: Window size for the kernel estimator.
  • $\phi_{k}^{(j)}$: Dynamic principal component filters, obtained as the Fourier coefficients of the dynamic eigenvectors.
  • $\varphi_{j}(\omega)$: Dynamic eigenvectors of the spectral density matrix $F(\omega)$.
  • $dpc_{t}^{(j)}$: Dynamic principal component scores.
  • $\mathbf{dpc}$: Matrix of dynamic principal component scores.
  • $\mathbf{\Phi}$: Matrix of dynamic principal component filters.
  • $\lambda_{j}(\omega)$: $j$-th dynamic eigenvalue of the spectral density matrix $F(\omega)$.
II-B Granger Causality test
Granger Causality (GC) is a statistical hypothesis test used to determine whether one time series can predict another. The concept, introduced by Granger (1969), is based on the principle that if the prediction of a time series $Y_{t}$ can be improved by incorporating past values of another time series $X_{t}$, then $X_{t}$ is said to ”Granger-cause” $Y_{t}$.
@@ II-B Granger Causality test
II-C Proposed Approach
In high-dimensional networks of time series, analyzing causal connectivity between specific pairs of signals can be both challenging and computationally intensive due to the complexity of modeling interactions across all nodes. The approach to be followed would be to model the whole material if there were no high dimensionality problems. However, modeling under high dimensionality requires fitting highly parameterized models to the entire network, leading to difficult-to-interpret estimates and an increased risk of overfitting. To overcome these challenges, we propose a practical and easily implementable method that focuses on isolating the nodes/channels of interest by effectively reducing the dimensionality of the network. Our approach utilizes spectral dynamic principal component analysis to capture the essential variability of the background network in a low-dimensional representation. By partialling out the influence of other nodes through sDPCA, we can isolate the COI and accurately assess the causal connectivity between them using conventional Granger causality tests. This method simplifies the analysis by avoiding the need to model the entire high-dimensional network and enhances interpretability when the primary interest is in the causal relationship between specific signals. The proposed approach is given in the Procedure outline below.
Input: High-dimensional Networks of Signals $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$
Output: Determination of causal relationship between $Y^{*}_{t}$ and $X^{*}_{t}$
1. Identify Nodes of Interest: NOI:$\{Y^{*}_{t},X^{*}_{t}\}$.
In practice, each node can also be a cluster of nodes that can be represented by sDPCA scores.
2. Apply sDPCA:
  1. i
    Estimate spectral density matrix $F_{\mathbf{Z}}(\omega)$ for $\mathbf{Z}_{t}=\{Z_{1,t},\ldots,Z_{n-2,t}\}$.
  2. ii
    Compute dynamic eigenvectors $\varphi_{j}(\omega)$ and
eigenvalues $\lambda_{j}(\omega)$.
  • iii
    Derive dynamic principal component filters:
[…]
  • iv
    Calculate dynamic principal component scores:
    $\text{dpc}_{t}^{(j)}=\sum_{k=-q}^{q}\phi_{k}^{(j)}\mathbf{Z}_{t-k}$
  • 3. Partial Out Influence of Other Nodes:
    $Y_{t}=Y^{*}_{t}-\hat{E}\big{(}Y^{*}_{t}\mid\text{dpc}_{t}^{(1)},\ldots,\text{dpc}_{t}^{(k)}\big{)}$
    $X_{t}=X^{*}_{t}-\hat{E}\big{(}X^{*}_{t}\mid\text{dpc}_{t}^{(1)},\ldots,\text{dpc}_{t}^{(k)}\big{)}$
    4. Fit Restricted/Unrestricted Models:
    Unrestricted Model:
    Restricted Model:
    $Y_{t}=\gamma_{0}+\sum_{i=1}^{p}\gamma_{i}Y_{t-i}+\eta_{t}$
    5. Perform Granger Causality Tests:
    1. i
      Test null hypothesis $H_{0}:\beta_{j}=0$
    for all $j$ in the equation for $Y_{t}$.
  • ii
    Compute F-statistic:
    $F=\frac{(RSS_{R}-RSS_{U})/q}{RSS_{U}/(T-K)}$
    where:
    • $RSS_{R}$: Residual sum of squares of
  • restricted model (without $X_{t-j}$ terms).
  • $RSS_{U}$: Residual sum of squares of
  • unrestricted model.
  • $q$: Number of restrictions (parameters set to zero). $K$: Total number of parameters in the unrestricted model. iii Compare F-statistic with critical value
  • from $F$-distribution $F_{q,T-K}$.
  • iv
    Reject $H_{0}$ if $F$ exceeds critical value; conclude $X_{t}$ Granger-causes $Y_{t}$. v Repeat steps for testing causality from $Y_{t}$ to $X_{t}$.
  • Procedure 1 Identifying Causal Connectivity in HDN
    III Simulation study
@@
    The aim of this simulation study is to evaluate the effectiveness of spectral dynamic principal component analysis (sDPCA) combined with Granger causality (GC) testing in identifying true causal relationships within a high-dimensional network of time series data, specifically focusing on EEG-like data. The study investigates how well the proposed method isolates the channels of interest (COI) from the influences of other channels and accurately assesses their causal interactions. We conduct two sets of simulations: one with linear influences and another with nonlinear influences.
    III-A Linear Scheme
    In the linear scheme of the simulation study, we generate a network of time series data with predefined causal relationships and linear influences from additional channels.
    III-A1 Generating Initial Channels
    We begin by generating $N=40$ initial time series channels, denoted by $\{Z_{i}(t)\}_{i=1}^{N}$, over time points $t=1,\ldots,T$. The channels are divided into $N/2=20$ causal pairs, where each pair consists of a ”cause” and an ”effect” channel.
    For each cause channel $Z_{i}(t)$, where $i=1,\ldots,N/2$, we model it as a weighted sum of latent processes representing different EEG frequency bands plus noise:
    $Z_{i}(t)=a_{i}\,\delta(t)+b_{i}\,\theta(t)+c_{i}\,\alpha(t)+d_{i}\,\beta(t)+e_{i}\,\gamma(t)+\varepsilon_{i}(t),$
    where:
@@ III-A1 Generating Initial Channels
    • $\delta(t)$, $\theta(t)$, $\alpha(t)$, $\beta(t)$, and $\gamma(t)$ are latent processes corresponding to the EEG frequency bands:
      • Delta (0.50.50.50.54444 Hz): $\delta(t)$,
      • Theta (44448888 Hz): $\theta(t)$,
      • Alpha (888812121212 Hz): $\alpha(t)$,
      • Beta (1212121230303030 Hz): $\beta(t)$,
      • Gamma ($>$ $30$ Hz): $\gamma(t)$.
      Each latent process is modeled as an autoregressive process of order 2 (AR(2)):
      $\nu(t)=\phi_{1\nu}\nu(t-1)+\phi_{2\nu}\nu(t-2)+\eta_{\nu}(t),\quad\nu\in\{\delta,\theta,\alpha,\beta,\gamma\},$
      where $\eta_{\nu}(t)$ is white noise. The AR coefficients can expressed in terms of $r_{\nu}$ and $\varrho_{\nu}$:
      $\phi_{1\nu}=2r_{\nu}\cos(\varrho_{\nu}),$
      $\phi_{2\nu}=-r_{\nu}^{2}.$
      where $r_{\nu}$ is the damping factor, chosen close to $1$ (e.g., $10.8\leq r_{\nu}<$ 1) to ensure stationarity and control the bandwidth of the spectral peak. And, $\varrho_{\nu}$ is the angular frequency (in radians per sample), related to the desired frequency $f_{\nu}$ (in Hz) by:
      $\varrho_{\nu}=2\pi\frac{f_{\nu}}{f_{s}},$
      with $f_{s}$ being the sampling frequency. By modeling the latent processes to correspond to specific EEG frequency bands, we aim to simulate EEG-like signals more realistically.
    • $a_{i}$, $b_{i}$, $c_{i}$, $d_{i}$, and $e_{i}$ are random weights drawn from a uniform distribution to ensure variability among channels.
    • $\varepsilon_{i}(t)$ is white noise added to each channel.
    For each effect channel $Z_{j}(t)$, where $j=i+N/2$ and $i=1,\ldots,N/2$, we generate it as an autoregressive process influenced by its corresponding cause channel:
    $Z_{j}(t)=\sum_{k=1}^{p}\phi_{k}Z_{i}(t-k)+\varepsilon_{j}(t),$
    where:
    • $p=2$ is the order of the autoregressive process.
    • $\phi_{k}$ are AR coefficients ensuring the stability of the process.
    • $\varepsilon_{j}(t)$ is white noise added to the effect channel.
    III-A2 Extending the Network
    We extend the network by generating additional $N_{\text{ext}}=472$ channels, resulting in a total of $N_{\text{expanded}}=512$ channels. These additional channels $\{Z_{i}(t)\}_{i=N+1}^{N_{\text{expanded}}}$ are generated similarly to the cause channels but without being part of the predefined causal pairs:
    $Z_{i}(t)=a_{i}\,\delta(t)+b_{i}\,\theta(t)+c_{i}\,\alpha(t)+d_{i}\,\beta(t)+e_{i}\,\gamma(t)+\varepsilon_{i}(t).$
    III-A3 Introducing Linear Influences
    To simulate the effect of a high-dimensional network on the initial channels, we introduce linear influences from a subset of the additional channels. We select a set of connected nodes $\mathcal{C}\subset\{N+1,\ldots,N_{\text{expanded}}\}$ of size nofcn (number of connected nodes).
    For each initial channel $Z_{i}(t)$, $i=1,\ldots,N$, we randomly select $k$ influencing nodes from $\mathcal{C}$ and assign influence weights $W_{ij}$ drawn from a uniform distribution.
    The linear influence on channel $i$ at time $t$ is defined as:
    $L_{i}(t)=\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t),$
    where $\mathcal{I}_{i}$ is the set of influencing nodes for channel $i$. For each initial channel $Z_{i}(t)$ and each influencing node $Z_{j}(t)\in\mathcal{I}_{i}$, we assign an influence weight $W_{ij}$ drawn from a uniform distribution centered around a specified influence weight $\omega$:
    $W_{ij}\sim\text{Uniform}\left(\omega-\frac{\omega}{16},\,\omega+\frac{\omega}{16}\right),\quad\text{for }j\in\mathcal{I}_{i}.$
    We then update the initial channels by adding the linear influence:
    $\tilde{Z}_{i}(t)=Z_{i}(t)+L_{i}(t).$
    The updated initial channels $\tilde{Z}_{i}(t)$ now contain both their original dynamics and the linear influence from the connected nodes.
    III-B Nonlinear Scheme
    In the nonlinear scheme, we introduce moderate nonlinearity into the influence of the connected nodes on the initial channels to study the robustness of the sDPCA and GC testing framework under nonlinear conditions.
    III-B1 Introducing Nonlinear Influences
    Similar to the linear simulation, we select a set of connected nodes $\mathcal{C}$ and assign influence weights $W_{ij}$ for each initial channel $Z_{i}(t)$.
    The linear influence is computed as before:
    $L_{i}(t)=\sum_{j\in\mathcal{I}_{i}}W_{ij}Z_{j}(t).$
    To introduce nonlinearity, we compute interaction terms between the influencing nodes for each initial channel:
    $Q_{i}(t)=\sum_{j\in\mathcal{I}_{i}}\sum_{k>j}W_{ij}W_{ik}Z_{j}(t)Z_{k}(t).$
    We apply a nonlinear function $f(\cdot)$ to the linear influence, such as the softplus function:
    $f(x)=\ln(1+e^{x}).$
    The total influence on channel $i$ at time $t$ is then given by:
    $S_{i}(t)=f\left(L_{i}(t)\right)+\omega Q_{i}(t),$
    where $\omega$ is a scaling parameter controlling the magnitude of the nonlinear interaction influence.
    We update the initial channels with the total influence:
    $\tilde{Z}_{i}(t)=Z_{i}(t)+S_{i}(t).$
    III-C Removing Influences
    III-C1 Removing Influences Using sDPCA
    To isolate the channels of interest from the influences of the connected nodes, we apply sDPCA to the time series data of the connected nodes $\{Z_{j}(t)\}_{j\in\mathcal{C}}$. We obtain principal component scores $\{Z_{\text{pc},k}(t)\}_{k=1}^{N_{\text{dpc}}}$, where $N_{\text{dpc}}$ is the number of dynamic principal components retained. For each adjusted initial channel $\tilde{Z}_{i}(t)$, we regress it on the principal component scores and their interaction terms:
    $\tilde{Z}_{i}(t)=\sum_{k=1}^{N_{\text{dpc}}}\varphi_{ik}Z_{\text{pc},k}(t)+\sum_{k\neq l}\psi_{ikl}Z_{\text{pc},k}(t)Z_{\text{pc},l}(t)+\epsilon_{i}(t).$
    We then use the residuals $\hat{\epsilon}_{i}(t)$ as the adjusted time series for further analysis:
    $\hat{Z}_{i}(t)=\hat{\epsilon}_{i}(t).$
    III-C2 Granger Causality Testing
    Using the adjusted time series $\{\hat{Z}_{i}(t)\}_{i=1}^{N}$, we fit vector autoregressive (VAR) models and perform Granger causality tests to identify causal relationships among the initial channels. The effectiveness of the sDPCA adjustment is evaluated by comparing the detected causal relationships with the ground truth.
    III-D Simulation Results
    To evaluate the performance of the proposed method, we compute accuracy metrics based on the detection of true causal relationships. Specifically, we compare the identified causal links with the known ground truth and compute measures such as accuracy rate.
    By conducting both linear and nonlinear simulations, we assess the robustness of the sDPCA combined with Granger causality testing in high-dimensional settings with complex interdependencies. The results provide insights into the method’s effectiveness in isolating the channels of interest and accurately detecting causal relationships, even in the presence of nonlinear influences.